OpenStudy (anonymous):

2(x-9)=46 what does x equal?

OpenStudy (anonymous):

at first divide both sides by 2 then you've got x-9 = 23 then add 9 on both side x-9+9= 23+9 so x = 32

OpenStudy (anonymous):

thanks so much! :D